Sound travels through water at an approximate rate of 1500 meters/second.
a)Write an equation describing d, the distance in meters that sound travels in t seconds.
-I know that d=rt so that would make the equation d=1500t
b)Graph the equation from part (a) by choosing three ordered pairs (t,d) that satisfy the equation. Label the coordinates of the 3 points you used. Be sure to label your axes.
-I'm not exactly sire how to graph this equation especially on such a small graph
(c) What is the slope of the line you graphed in part (b.)
(d) A whale-watching boat's sonar determines that it takes sound waves a total of three seconds to reach a pod of whales and return to the boat. How many metera is the pod of whales from the boat?

Answer by stanbon(75887)

a) Yes. d=1500t
b) If t=0 , d=0 gives plotting point (0,0)
If t=1 , d=1500 gives plotting point (1,1500)
The horizontal axis is the "t" axis; The vertical is the "d" axis.
c) Between those two points the change in d is 1500-0 = 1500
the change in t is 1-0 = 1. Slope = 1500/1= 1500
d) In 3 seconds the wave travels 3(1500)meters = 4500 meters.
Let the distance to the pod be x. The round trip for the wave
is 2x. Therefore 2x=4500 meters. x = 2250 meters.
Distance to the pod is 2250 meters.
Cheers,
Stan H.